# Stress-stress correlations in two-dimensional amorphous and crystalline solids

### v1

@@ I Introduction
Despite the success of VCT in describing the emergent (static) elasticity of amorphous solids, several open questions remain. In particular, the similarities with elasticity theory are in fact striking, in particular when ET is considered in its stress-only formulation. Is VCT really needed to derive the static stress-stress correlations in amorphous solids? Which would be the predictions of vanilla elasticity theory and how would they differ from those arising from VCT?
@@
The purpose of this work is to revisit the question of static stress-stress correlations in 2D amorphous and crystalline systems. As we will explicitly demonstrate and corroborate with experimental and simulation data, the VCT predictions for the static stress-stress correlations are identical to those of vanilla elasticity theory for isotropic crystals. In other words, the static stress response of amorphous solids appears of the same form of that of isotropic crystals, including the presence of pinch-point singularities that were previously ascribed as exclusive characteristic of amorphous solids.
@@ II Stress-stress correlations from vanilla elasticity theory
The ensemble average in Eq. (12) refers to the thermodynamic ensemble that is defined by the Boltzmann probability distribution $e^{-\frac{S}{k_{B}T}}$. In other words, in thermal systems, $g\equiv k_{B}T$. Since the value of $g$ affects our final results only up to an overall constant factor, for simplicity, we will set $g=1$ in the rest of the manuscript. Before proceeding, we notice that the choice of the ensemble over which the average is performed is the main difference between the computation from elasticity theory (sketched in this section) and the analogous computation for athermal granular systems presented in [6]. There, the ensemble average is not on the Boltzmann thermal distribution and $g\neq k_{B}T$. In that case, the average is performed using an equiprobability ansatz that is the basis of the Edwards stress ensemble (see [18] for a recent review). As we will concretely see, this difference will not alter the form of the correlations as a function of the wave-vector q nor the angle $\theta$.
@@ II.2 Stress-Stress Correlations for a 2D Square Lattice
Figure 1: Pair distribution function $g(\tilde{r})$: (TLS) triangular simulation, (SLS) square simulation, (AE) amorphous experiment, and (TLE) triangular experiment. $\tilde{r}$ denotes the radial length-scale normalized by the nearest neighbor distance for crystalline systems and the average distance between nearby particles for the amorphous packing. Inset images provide snapshots of the corresponding structures.
@@ III.1 Experimental details
In the experiments, we apply photoelastic techniques to measure the inter-particle forces. Initially, we adjust the structure arrangement of the crystals and grain boundaries to achieve zero pressure. Subsequently, isotropic quasi-static compression is applied via boundary displacement, with a maximum compressive ratio of 1.04%. Following this protocol, we obtain 8 independent samples. For further information of the experimental apparatus and the photoelastic techniques, please refer to Refs. [19, 20].
@@ III.2 Simulation details
Figure 3: Normalized stress fluctuation correlations $C_{ijkl}(\theta)$ obtained from the amorphous bidisperse experimental packing (AE in Fig. 1). The data are obtained by integrating the stress fluctuation correlations $C_{ijkl}$ over $q$ to retain only the angular dependence, where $q$ is the magnitude of the Fourier space $q$-vector. The results are the ensemble average over $10$ independent experimental datasets and the error bars indicate the corresponding statistical errors. The solids lines are the theoretical predictions from Eq. 20.
@@ IV Amorphous packing versus isotropic crystal
In Fig. 3, we show the stress fluctuation correlations $C_{ijkl}(\theta)$ as a function of the angular variable $\theta\in[0,2\pi]$ for the same 2D experimental amorphous packing. To obtain the $\theta$ dependence of $C_{ijkl}$, we integrate $C_{ijkl}(q,\theta)$ over $q$, where $q$ is the magnitude of the Fourier space $q$-vector. The blue symbols indicate the experimental results averaged over $10$ independent datasets. The error bars indicate the corresponding statistical errors. In the same figure, the solid red lines show the theoretical predictions from 2D elasticity theory, Eq. 20. The agreement between the experimental data and the theoretical predictions is evident. We notice that the theoretical predictions are based solely on elasticity theory. Despite this might seem surprising, we recall that, in the regime in which only quadrupole screening is relevant in amorphous solids, the mechanical response of athermal amorphous systems is well described by elasticity theory, albeit with renormalized and parameter dependent elastic moduli. It is therefore not surprising that elasticity theory correctly captures the form of the correlations. On the other hand, whenever dipole screening becomes relevant, our formulae (and also the identical ones derived from the original VCT theory) will fail, since they will not be able to capture the emergent plastic screening length-scale that needs to be incorporate into the elasticity framework as done in [27].
Figure 4: 2D stress fluctuation correlations ($C_{xxxx}$, $C_{xxyy}$, $C_{xxxy}$, $C_{yyyy}$, $C_{yyxy}$, and $C_{xyxy}$) as a function of $\theta$, obtained from triangular lattice simulation (TLE in Fig. 1). The data are normalized by the maximum value of $C_{xxxx}$.
@@
Figure 5: Normalized stress fluctuation correlations ($C_{xxxx}$, $C_{xxyy}$, $C_{xxxy}$, $C_{yyyy}$, $C_{yyxy}$, and $C_{xyxy}$) as a function of $\theta$, obtained from the triangular lattice simulation (TLS in Fig. 1). The normalization is performed by dividing each component by the maximum value of $C_{xxxx}$. The solid red lines represent the theoretical predictions in Eq. 20. The error bars denote statistical error in 20 independent simulation samples.
@@
Figure 6: Normalized 2D correlation functions obtained from square lattice simulation. 20 groups of independent simulation are concerned.
@@ V Stress-stress correlations in anisotropic crystals
Figure 7: Normalized correlation functions obtained from square lattice simulation (SLS in Fig. 1). The solids lines correspond to the theoretical predictions, Eq. 24. The error bars denote statistical error in 20 independent simulation samples.
@@ VI Discussion
In this work, we revisited the problem of static stress-stress correlations in 2D solids with a particular emphasis on the contrast between amorphous and crystalline structures. After presenting a general derivation of these correlations from vanilla elasticity theory, we have focus on two benchmark cases: (i) isotropic 2D solid (triangular crystalline structure) and (ii) anisotropic 2D solid (square lattice).
@@ Appendix A Isotropic 2D crystal: experiments
Figure 8: Normalized stress fluctuation correlations $C_{ijkl}(\theta)$ obtained from the isotropic crystalline experimental packing (TLE in Fig. 1). The results are the ensemble average over $7$ independent experimental datasets and the error bars indicate the corresponding statistical errors. The solids lines are the theoretical predictions from Eq. 20.

### v2

@@ I Introduction
Despite the success of VCT in describing the emergent (static) elasticity of amorphous solids, several open questions remain. In particular, the similarities with elasticity theory are in fact striking, in particular when ET is considered in its stress-only formulation. Is VCT really needed to derive the static stress-stress correlations in amorphous solids? Which would be the predictions of vanilla elasticity theory [1, 2] and how would they differ from those arising from VCT?
@@
The purpose of this work is to revisit the question of static stress-stress correlations in 2D amorphous and crystalline systems. As we will explicitly demonstrate and corroborate with experimental and simulation data, the VCT predictions for the static stress-stress correlations are identical to those of vanilla elasticity theory for isotropic crystals [1, 2]. In other words, the static stress response of amorphous solids appears of the same form of that of isotropic crystals, including the presence of pinch-point singularities that were previously ascribed as exclusive characteristic of amorphous solids (see, e.g., [6] and references therein).
@@ II Stress-stress correlations from vanilla elasticity theory
The ensemble average in Eq. (12) refers to the thermodynamic ensemble that is defined by the Boltzmann probability distribution $e^{-\frac{S}{k_{B}T}}$. In other words, in thermal systems, $g\equiv k_{B}T$. Since the value of $g$ affects our final results only up to an overall constant factor, for simplicity, we will set $g=1$ in the rest of the manuscript. Before proceeding, we notice that the choice of the ensemble over which the average is performed is the main difference between the computation from elasticity theory (sketched in this section) and the analogous computation for athermal granular systems presented in [6]. There, the ensemble average is not on the Boltzmann thermal distribution and $g\neq k_{B}T$. In that case, the average is performed using an equiprobability ansatz that is the basis of the Edwards stress ensemble (see [18] for a recent review). The Edwards ensemble is based on the fact that volume is conserved and compactivity $\chi_{V}$ equilibrates, as temperature does in the Boltzmann ensemble. One can then compute the density of jammed states $\Omega(V)$ as a function of the volume $V$ and, finally, build the Edwards entropy, the conjugate variable of the compactivity $\chi_{V}$. This framework is particularly useful for jammed granular matter [19], and it has been recently experimentally validated in three-dimensional granular systems [20]. As we will concretely see, this difference will not alter the form of the correlations as a function of the wave-vector q nor the angle $\theta$.
@@ II.2 Stress-Stress Correlations for a 2D Square Lattice
Before continuing, we clarify that our work does not introduce new theoretical models but rather relies on established results from standard elasticity theory. For athermal amorphous solids, more sophisticated theories, able to derive also an “equation of state”, have been already developed [21, 22]. Numerical studies of stress-stress correlations in deeply supercooled liquids also exist, e.g., [23].
Figure 1: Pair distribution function $g(\tilde{r})$: (TLS) triangular simulation, (SLS) square simulation, (AE) amorphous experiment, and (TLE) triangular experiment. $\tilde{r}$ denotes the radial length-scale normalized by the nearest neighbor distance for crystalline systems and the average distance between nearby particles for the amorphous packing. The images below provide snapshots of the corresponding structures.
@@ III.1 Experimental details
In the experiments, we apply photoelastic techniques to measure the inter-particle forces. Initially, we adjust the structure arrangement of the crystals and grain boundaries to achieve zero pressure. Subsequently, isotropic quasi-static compression is applied via boundary displacement, with a maximum compressive ratio of 1.04%. Following this protocol, we obtain 7 independent samples. For further information of the experimental apparatus and the photoelastic techniques, please refer to Refs. [24, 25].
@@ III.2 Simulation details
Figure 3: Normalized stress fluctuation correlations $C_{ijkl}(\theta)$ obtained from the amorphous bidisperse experimental packing (AE in Fig. 1). The data are obtained by integrating the stress fluctuation correlations $C_{ijkl}$ over $q$ to retain only the angular dependence, where $q$ is the magnitude of the Fourier space $q$-vector. The results are the ensemble average over $10$ independent experimental datasets and the error bars indicate the corresponding statistical errors. The solids lines are the theoretical predictions from Eq. (20).
@@ IV Amorphous packing versus isotropic crystal
In Fig. 3, we show the stress fluctuation correlations $C_{ijkl}(\theta)$ as a function of the angular variable $\theta\in[0,2\pi]$ for the same 2D experimental amorphous packing. To obtain the $\theta$ dependence of $C_{ijkl}$, we integrate $C_{ijkl}(q,\theta)$ over $q$, where $q$ is the magnitude of the Fourier space $q$-vector. The blue symbols indicate the experimental results averaged over $10$ independent datasets. The error bars indicate the corresponding statistical errors. In the same figure, the solid red lines show the theoretical predictions from 2D elasticity theory, Eq. (20). The agreement between the experimental data and the theoretical predictions is evident. We notice that the theoretical predictions are based solely on elasticity theory. Despite this might seem surprising, we recall that, in the regime in which only quadrupole screening is relevant in amorphous solids, the mechanical response of athermal amorphous systems is well described by elasticity theory, albeit with renormalized and parameter dependent elastic moduli. It is therefore not surprising that elasticity theory correctly captures the form of the correlations. On the other hand, whenever dipole screening becomes relevant, our formulae (and also the identical ones derived from the original VCT theory) will fail, since they will not be able to capture the emergent plastic screening length-scale that needs to be incorporate into the elasticity framework as done in [32]. In the context of athermal granular matter, the regime in which dipole screening is negligible corresponds to the high-pressure limit (see, e.g., Fig.3 in [33]).
Figure 4: 2D stress fluctuation correlations ($C_{xxxx}$, $C_{xxyy}$, $C_{xxxy}$, $C_{yyyy}$, $C_{yyxy}$, and $C_{xyxy}$) as a function of $\theta$, obtained from triangular lattice simulation (TLE in Fig. 1). The data are normalized by the maximum value of $C_{xxxx}$. Pinch-point singularities in the $q\rightarrow 0$ limit emerge in the correlation functions, since the latter are only a function of the angle $\theta$ in the continuous small $q$ regime.
@@
Figure 5: Normalized stress fluctuation correlations ($C_{xxxx}$, $C_{xxyy}$, $C_{xxxy}$, $C_{yyyy}$, $C_{yyxy}$, and $C_{xyxy}$) as a function of $\theta$, obtained from the triangular lattice simulation (TLS in Fig. 1). The normalization is performed by dividing each component by the maximum value of $C_{xxxx}$. The solid red lines represent the theoretical predictions in Eq. (20). The error bars denote statistical error in 20 independent simulation samples.
@@
Figure 6: Normalized 2D correlation functions obtained from square lattice simulation. To improve statistics, $20$ groups of independent simulation are concerned. Pinch-point singularities in the $q\rightarrow 0$ limit are evident in all correlation functions.
We emphasize that our theoretical framework is based on continuum elasticity theory, which is valid only in the long-wavelength limit, i.e., for wave vectors $q$ much smaller than the inverse lattice spacing $a$. To substantiate this point, Fig.7 presents an analysis of the stress fluctuation correlation function $C_{xyxy}$ evaluated at different fixed values of $q$. In panel (a), we show $C_{xyxy}$ in the two-dimensional $(q_{x},q_{y})$ plane, with concentric circles indicating the $q$-cuts plotted in panel (b). As clearly seen in Fig.7(b), the theoretical prediction from Eq. (20) matches the simulation results well at low $qa$. However, at higher $qa$, significant deviations appear, indicating the breakdown of the continuum approximation at shorter length scales.
Figure 7: (a) 2D stress fluctuation correlation function $C_{xyxy}$ as a function of the normalized wave-vector $\textbf{q}a$ from the triangular lattice simulation. The various circles indicate the $q$-cuts shown in panel (b). (b) Angular dependence of the same correlation function at constant $q$. The dashed black line is the theoretical prediction from elasticity theory, Eq. (20) in the main text. The shaded regions are the statistical errors associated to the numerical data.
@@ V Stress-stress correlations in anisotropic crystals
Figure 8: Normalized correlation functions obtained from square lattice simulation (SLS in Fig. 1). The solids lines correspond to the theoretical predictions, Eq. (24). The error bars denote statistical error in 20 independent simulation samples.
@@ VI Discussion
In this work, we revisited the problem of static stress-stress correlations in 2D solids with a particular emphasis on the contrast between amorphous and crystalline structures. After presenting a general derivation of these correlations from vanilla elasticity theory [1, 2], we have focus on two benchmark cases: (i) isotropic 2D solid (triangular crystalline structure) and (ii) anisotropic 2D solid (square lattice).
@@ Appendix A Isotropic 2D crystal: experiments
Figure 9: Normalized stress fluctuation correlations $C_{ijkl}(\theta)$ obtained from the isotropic crystalline experimental packing (TLE in Fig. 1). The results are the ensemble average over $7$ independent experimental datasets and the error bars indicate the corresponding statistical errors. The solids lines are the theoretical predictions from Eq. (20).